(a) Find the domain and range of the function f(x) = log(8x).
(b) Find the inverse of f(x) = log(x).
f -1(x) =

Answer by stanbon(75887)

(a) Find the domain and range of the function f(x) = log(8x).
----
Domain?
8x must be >0, so x must be > 0
-----------
Domain: All real Numbers greater than zero.
==================================================

(b) Find the inverse of f(x) = log(x).
Interchange x and y to get:
x = log(y)
Solve for "y":
y = 10^x
--------------
f -1(x) = 10^x
==================================================
Cheers,
Stan H.
================